Find three consecutive even integers such that the difference of the third and half of the first is 14
Answer by CubeyThePenguin(3113)

consecutive even integers: (x-2), x, (x+2)

(x+2) - (x-2)/2 = 14
x + 2 - (x/2 - 1) = 14
x + 2 - x/2 + 1 = 14
x/2 = 11
x = 22

The integers are 20, 22, and 24.